OpenStudy (anonymous):

use the formula P=Ie^kt. A bacterial culture has an initial population of 500. If its population grows to 7000 in 4hrs, what will it be at the end of 6 hrs?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

at t=0, P=500 use that to find A 500=Ae^k*0 A=500 Then find K P=500e^k*t 7000=500e^k*4 after that, put k back in and solve for P using t=6

OpenStudy (anonymous):

i still dont get it, maybe im putting it in the calculator wrong but thanks for your help